OpenStudy (anonymous):

Can someone please help me write the standard form of the quadratic function that has the indicated vertex and whose graph passes through the given point. Vertex: (-5/2,0) Point: (-7/2,-16/3)

OpenStudy (campbell_st):

well the standard vertex form is \[y = a(x -h)^2 + k\] where the vertex is (h, k) so you can substitute the vertex directly into the equation. next to find a, after substituting the vertex, substitute the point into the equation to get the value of a. lastly, to get the equation in standard form with the vertex substituting and having the value of a, expand the perfect square and distribute a then collect like terms.. hope this helps

OpenStudy (anonymous):

would just substituting the vertex be y=a(x-4)^2-1? @campbell_st

OpenStudy (anonymous):

how do i substitute the point if there are two numbers for the point?

OpenStudy (campbell_st):

nope the vertex is (-5/2, 0) so h = -5/2 and k = 0

OpenStudy (campbell_st):

and if you have a point (x, y) and an equation y =a(x -h)^2 + k you have an x and y value in the equation, replace then with the values from the point...

OpenStudy (anonymous):

sorry lol i was looking at a different problem so now i have -16/3=a(7/2-(-5/2)^2+0 is that correct? @campbell_st

OpenStudy (campbell_st):

thats correct, so solve for a then you'll have your equation

OpenStudy (anonymous):

what do i need to do to solve for a??

OpenStudy (campbell_st):

well simplify -16/3 = a(7/2 - (-5/2))^2 or -16/3 = a(72/ + 5/2)^2 -16/3 = a(12/2)^2 etc...

OpenStudy (anonymous):

so then a would be -4/27?

OpenStudy (anonymous):

so the standard form with the given information would be -16/3=-4/27(7/2-(-5/2)^2?

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

this is what I have, how would I write this in the proper standard form?

OpenStudy (campbell_st):

ok..so you have the value of a then \[y = -\frac{4}{27}(x - \frac{5}{2})^2\] is the vertex form... distribute to get the standard form so its \[y =- \frac{4}{27}(x^2 - 5x + \frac{25}{4})\] just distribute for the final standard form...

OpenStudy (anonymous):

so would the final standard form be y=-(2x-5)^2/27? i dont know if i did this right

OpenStudy (campbell_st):

well in standard form, I'd expect you'd need to distribute the -4/27

OpenStudy (anonymous):

when i try distrbuting the -4/27 i keep getting -(2x-5)^2/27

OpenStudy (campbell_st):

well my starting point is \[y = -\frac{4}{27}(x^2 - 5x + \frac{25}{4}) = \frac{-4x^2 + 20x -25}{27}\] just as a guess...

OpenStudy (anonymous):

well then wouldnt it be -4x^2+20x-100/27?